THE WEIGHT OF AN OBJECT ON THE MOON VARIES DIRECTLY AS ITS WEIGHT ON EARTH. WITH ALL OF HIS GEAR ON, NEIL ARMSTRONG WEIGHED 360 POUNDS ON EARTH. WHEN HE BECAME THE FIRST PERSON TO STEP ON THE MOON ON JULY 20, 1969, HE WEIGHED 60 POUNDS. TARA WEIGHS 108 POUNDS ON EARTH. WHAT WOULD SHE WEIGH ON THE MOON. Answer by stanbon(75887) 
 THE WEIGHT OF AN OBJECT ON THE MOON VARIES DIRECTLY AS ITS WEIGHT ON EARTH.
---
m = ke
------
Solve for "k" using " WITH ALL OF HIS GEAR ON, NEIL ARMSTRONG WEIGHED 360 POUNDS ON EARTH. WHEN HE BECAME THE FIRST PERSON TO STEP ON THE MOON ON JULY 20, 1969, HE WEIGHED 60 POUNDS.
60 = k*360
k = 1/6
---
Equation:
m = (1/6)e
----
TARA WEIGHS 108 POUNDS ON EARTH. WHAT WOULD SHE WEIGH ON THE MOON.
m = (1/6)108
m = 18 lbs (Tara's weight on the moon)
===================
Cheers,
Stan H.
===============